determine the average gravitational force between a 1.05 solar mass and a planet whose orbital speed is 20,000 m/s if its distance from the star is 149 million kilometers at its nearest and 150 million kilometers at its farthest.

1
Expert's answer
2022-01-16T13:08:57-0500

Given quantities:

"M_1=1.05M_s"

"v = 20000 \frac{m}{s}"

"R = 150 *10^9 m"

"F-?"


"F = G*\frac{M_1*M_2}{R^2}" ="G*\frac{1.05M_s^2}{R^2}=\frac{6.67*10^{-11}*4^{60}}{225*10^{20}}=1.2*10^{28}N"